find four consecutive even integers such that the sum of the second and fourth is 76?
Answer #1
X, X+2, X+4, X+6 X+2+X+6=76 2X+8=76 2X=76-8 2X=68 X=68/2 X=34 ANSWER FOR THE SMALLEST NUMBER. THE SECOND=34+2=36 THE FOURTH=34+6=40 PROOF 36+40=76 76=76
answered by: jhenssine